# PertEval: Unveiling Real Knowledge Capacity of LLMs with Knowledge-Invariant Perturbations

### v1

@@ Appendix A Related Work
A.1 LLMs’ Knowledge Capacity Evaluation Benchmarks
@@ A.2 Trustworthiness of Knowledge Capacity Evaluation
Appendix B Discussion about PertEval
The ultimate goal of PertEval is not only to unveil the real knowledge capacity of LLMs, but also to guide the refinement of LLMs in enhancing knowledge capacity. To this end, we discuss potential methods and future directions for enhancing LLMs’ knowledge capacity with PertEval:
Supervised fine-tuning with perturbed data. PertEval can be utilized to generate various forms of multiple-choice question data. For knowledge that requires memorization such as history and laws, new data generated by PertEval can be used to enhance the ability of LLMs to adjust to various complicated test scenarios. Indeed, we conduct a primary experiment using llama-3-8b-instruct (see Appendix D.4) and reveal the feasibility of this solution.
Extending PertEval to open-ended test data. The current PertEval adapts only to multiple-choice questions considering their precision of measurement and cost effectiveness. To further evaluate the knowledge capacity of LLMs in all kinds of domains, a future direction is to extend PertEval to open-ended data types such as open question answering. Potential challenges in this direction include but not limited to how to define the correctness of output answer given perturbations and how to ensure knowledge invariance in complicated questions such as solving math application questions.
Appendix C Details of Methodology
C.1 Content-level Perturbation: Knowledge-invariant Paraphrasing
@@
Figure 4: Prompt template for the rewriter LLM.
@@
C.2 Format-level Perturbation: Question Format Refactoring
@@ A $x=1$; B $x=2$; C $x=3$; D $x=4$
Appendix D Details of Experiments
@@
D.1 Checking the Knowledge Invariance of Perturbation Strategies
In this part, we utilize the proposed knowledge invariance checking methods to measure the validity of perturbations in terms of knowledge invariance. The prompt template for the LLM-based knowledge invariance scoring is presented in Figure 5 and 6. Results of LLM-based knowledge invariance scoring are shown in Figure 1. Results of overall performance stability testing on mastered questions are presented in Figure 3. We obtain several conclusions from these experimental results, as given in the following:
@@ D.1 Checking the Knowledge Invariance of Perturbation Strategies
D.2 Probing LLMs’ Consistent Knowledge Capacity
@@
$E[\text{ACC@Consist}]=E[X/n]=\frac{1}{n}E[X]=1/k^{2}$ (4)
@@
D.3 Response Pattern Analysis for LLMs in PertEval
In this part, we analyze response pattern for LLMs in the original and perturbed data, which can provide us a deep insight into the influence of knowledge-invariant perturbations on LLMs’ question answering procedure. Specifically, response patterns consist of correct choice, invalid choice, extra multiple choice, wrong single choice and wrong multiple choice. Their definition are listed in the following:
  • Correct choice: The LLM selects and only selects all correct options. The ratio of correct choices is equvalent to micro accuracy.
  • Invalid choice: The LLM selects none of options. This could be caused by the LLM’s wrong reasoning process or invalid output format.
  • Extra multiple choice: The LLM selects not only all correct options, but also extra incorrect options. This means that the LLM fails to filter out incorrect options from the option list.
  • Wrong single choice: The LLM selects only a single option, and the option is wrong. This means that the LLM fails to recognize the correct option in the option list.
  • Wrong multiple choice: The LLM selects multiple options which do not cover all correct options.
Given the response pattern taxonomy, we counted the ratio of response patterns in original and perturbed data for all selected LLMs, as shown in Figure 3, 8, 9, 10, 11 and 12.
(a) Original data
(b) Perturbed data with composite perturbation
Figure 8: Response patterns of gpt-3.5-turbo in all datasets.
(a) Original data
(b) Perturbed data with composite perturbation
Figure 9: Response patterns of glm-3-turbo in all datasets.
(a) Original data
(b) Perturbed data with composite perturbation
Figure 10: Response patterns of gemini-1.0-pro in all datasets.
(a) Original data
(b) Perturbed data with composite perturbation
Figure 11: Response patterns of llama-3-8b-instruct in all datasets.
(a) Original data
(b) Perturbed data with composite perturbation
Figure 12: Response patterns of mistral-7b-instruct-v0.2 in all datasets.
D.4 How to Enhance LLMs’ Knowledge Capacity Using PertEval?
@@
Table 10: PDR of LLaMA-8B-Instruct before and after finetuning on perturbed datasets. Underlined datasets are used for finetuning. CT denotes ChangeType. KP denotes KnInvPara. SP denotes SwapPos. For example, F(CT) means finetuning LLaMA-8B-Instruct on world history and professional medicine datasets perturbed by ChangeType.

### v2

@@ Appendix A Related Work
This section complements related works especially in LLM evaluation for Section 2.
A.1 Benchmarks for Knowledge Capacity Evaluation
@@ A.2 Trustworthiness of Knowledge Capacity Evaluation
Appendix B Details of Methodology
B.1 Content-level Perturbation: Knowledge-invariant Paraphrasing
@@
Figure 5: Prompt template for the rewriter LLM. The expected similarity score is used to control parphrasing of the rewriter LLM, which is set to 0.6 in all experiments.
@@
B.2 Format-level Perturbation: Question Format Refactoring
@@ A $x=1$; B $x=2$; C $x=3$; D $x=4$
B.3 Correct Response Consistency
Similar to the confusion matrix, terms in ROP can be written in a response consistency matrix, as shown in Table 10. Here “$C$”/“$I$” in the first position denotes “consistently”/“inconsistently”. “$C$”/“$W$” in the second position denotes “correct”/“wrong”. For instance, $CC$ denotes consistently correct.
Table 10: Response transition matrix.
Number Correct after perturbation $\sigma$ Wrong after perturbation $\sigma$
Correct before perturbation $\sigma$ $CC$ $IC$
Wrong before perturbation $\sigma$ $IW$ $CW$
Appendix C Experiment Setup
C.1 Details of Dataset
@@
C.2 Evaluation Setting
Model Setting. For close-sourced LLMs (gpt-4-turbo, gpt-3.5-turbo, glm-3-turbo, gemini-1.0-pro), we use their APIs for evaluation. The temperature parameter of each model is set to 0.2 to reduce output uncertainty. For open-sourced LLMs (llama-3-8b-instruct, mistral-7b-instruct-v0.2), we locally deploy their huggingface version to our own server and utilize default settings for evaluation.
Environment Setting. All experiments are completed on a Linux server with Intel(R) Xeon(R) Platinum 8269CY CPUs @ 2.50GHz and one NVIDIA A100 GPU (40G). GPUs are used for deploying and fine-tuning open-sourced models. The version of Python is 3.9. The version of the torch package is 2.2.0. The version of the transformers package is 4.39.3.
Prompt Setting. In our experiments, we use prompt templates to generate prompts of questions for evaluation. Examples using prompt templates are shown in Figure 8 and 9. After getting model outputs, we transform them to json data for further analyses and utilize regular expressions to enhance the robustness of transformation. Details are available at our code.
Supervised Fine-tuning Setting. In Appendix D.3, we fine-tune llama-3-8b-instruct using PertEval-generated data to explore the utility of PertEval in enhancing the knowledge capacity of LLMs. The fine-tuning process is done using LoRA [65] on single GPU implemented in LLaMA-Factory [66]. The batch size is 1, since the number of training data is 237 (High School World History) + 272 (Professional Medicine) = 479. The learning rate is 5e-5. The number of training epoch is 3.
Please select the correct option(s) from the following options given the question: Question: Let k be the number of real solutions of the equation $e^{x}+x-2=0$ in the interval [0, 1], and let n be the number of real solutions that are not in [0, 1]. Which of the following is true? Options: A k = 0 and n = 1 B k = 1 and n = 0 C k = n = 1 D k > 1 Your output must strictly follow this format: {"answer": <the list of selected options, e.g., ["A", "B", "C", "D"]>} Your output:
Figure 8: Prompt template for multiple-choice question answering.
Please judge whether each of the options is correct or incorrect given the question: Question: Let k be the number of real solutions of the equation $e^{x}+x-2=0$ in the interval [0, 1], and let n be the number of real solutions that are not in [0, 1]. Which of the following is true? Options: A k = 0 and n = 1 B k = 1 and n = 0 C k = n = 1 D k > 1 Your output must strictly follow this format: {"A": <"True" or "False">, "B": <"True" or "False">, "C": <"True" or "False">, "D": <"True" or "False">} Your output:
Figure 9: Prompt template for multiple-judgement question answering (ChangeType).
@@ C.2 Evaluation Setting
C.3 Experiment Cost
Costs in # Tokens. Overall, we conduct the following experiment steps that require token costs:
  1. E0.
    Generating PertEval benchmark.
  2. E1.
    Knowledge-invariance validation - obtain knowledge invariance score.
  3. E2.
    Main experiment - obtain ACC@Consist.
  4. E3.
    Fine-grained perturbation analysis - obtain ROP and PDR.
Exact costs w.r.t. experiments are presented as follows:
Table 12: Exact token costs w.r.t. experiments.
Index # Input tokens # Output tokens
E0 3,542,896 374,967
E1 776,353 101,675
E2 3,248,124 190,476
E3 9,343,200 336,996
Costs in USD. For proprietary LLMs, the cost is from the pricing strategy of LLM APIs (proportional to # input/output tokens). For open-sourced LLMs, the cost originates from the pricing strategy of servers for LLM deployment (proportional to the running time of servers). We count overall costs in USD per experiment, as shown in Table 13.
Table 13: Costs in USD per experiment.
Exp Index Price (USD)
E0 41.79
E1 10.81
E2 13.09
E3 51.47
Total 117.16
Appendix D Verifying the Knowledge Invariance of Perturbation Strategies
In this part, we utilize the proposed knowledge invariance verification methods to measure the validity of perturbations in terms of knowledge invariance. The prompt template for the LLM-based knowledge invariance scoring is presented in Figure 6 and 7. Results of LLM-based knowledge invariance scoring are shown in Figure 2. Results of overall performance stability testing on mastered questions are presented in Figure 4. We obtain several conclusions from these experimental results, as given in the following:
@@ Appendix D Verifying the Knowledge Invariance of Perturbation Strategies
D.1 Probing LLMs’ Consistent Knowledge Capacity
@@
$E[\text{ACC@Consist}]=E[Z/n]=\frac{1}{n}E[Z]=1/k^{2}$ (4)
@@
D.2 Response Pattern Analysis for LLMs in PertEval
In this part, we analyze response pattern for LLMs in the original and perturbed data, which can provide us a deep insight into the influence of knowledge-invariant perturbations on LLMs’ question answering procedure. The definition of response patterns are listed in the following:
  • Single Correct Selection: The LLM selects and only selects the correct option. The ratio of correct choices is equvalent to micro accuracy.
  • Single Wrong Selection: The LLM selects only a single option, and the option is wrong. This means that the LLM fails to recognize the correct option in the option list.
  • Invalid Selection: The LLM selects none of options. This could be caused by the LLM’s wrong reasoning process or invalid output format.
  • Multiple Selection Including the Correct One: The LLM selects not only the correct option, but also extra incorrect options. This means that the LLM fails to filter out incorrect options from the option list.
  • Multiple Selection w/o the Correct One: The LLM selects multiple options which do not cover the correct option.
Given the response pattern taxonomy, we counted the ratio of response patterns in original and perturbed data for all selected LLMs, as shown in Figure 4, 11, 12, 13, 14 and 15.
Figure 11: Response patterns of gpt-3.5-turbo in all datasets. Left: original data. Right: perturbed data.
Figure 12: Response patterns of glm-3-turbo in all datasets. Left: original data. Right: perturbed data.
Figure 13: Response patterns of gemini-1.0-pro in all datasets. Left: original data. Right: perturbed data.
Figure 14: Response patterns of llama-3-8b-instruct in all datasets. Left: original data. Right: perturbed data.
Figure 15: Response patterns of mistral-7b-instruct-v0.2 in all datasets. Left: original data. Right: perturbed data.
D.3 How to Enhance LLMs’ Knowledge Capacity Using PertEval?
@@
Table 14: PDR of LLaMA-8B-Instruct before and after fine-tuning on perturbed datasets. Underlined datasets are used for fine-tuning. CT denotes ChangeType. KP denotes KnInvPara. SP denotes SwapPos. For example, F(CT) means fine-tuning LLaMA-8B-Instruct on world history and professional medicine datasets perturbed by ChangeType.
@@
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 5.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [No] Our work mainly focuses on techniques for LLM evaluation. We proposed a novel evaluation toolkit for enhancing existing evaluation benchmark rather proprosing than a new dataset. Therefore, there is no risk for this work to have negative societal impacts caused by online data collection, such as the invasion of users’ privacy.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [Yes]
    2. (b)
      Did you include complete proofs of all theoretical results? [Yes] Due to page limitation, proofs are available at Appendix D.1.
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] The URL to our code and data is available at abstract.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Due to page limitation, details are available at Appendix C.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] It is too costly and timely expensive to conduct experiments for multiple times given the huge number of settings (assuming there are 6 models, 6 perturbations, 4 datasets and 10 running times, there are at least 6*6*4*10 = 1,440 groups of experiments). Substitutionally, in experiments, we set the temperature parameter to all LLMs as a relative small value ($\tau=0.2$) to reduce the uncertainty of LLMs’ outputs. We also open-sourced our code to ensure reproducibility.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] A description of computational resources is available at Appendix C.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We used LLMs (both APIs of close-sourced LLMs and weights of open-sourced LLMs) and open-sourced datasets in our work.
    2. (b)
      Did you mention the license of the assets? [No]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] The URL of the code and data of our work is available at abstract.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A] We did not collect private data from people. Instead, all APIs, model weights and evaluation data used in this work, are published online and accessible to everyone.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A] Data we used in our work is about knowledge capacity evaluation, and does not contain personally identifiable information.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]